OpenStudy (anonymous):

What's the implest form in x5y3z3/xy-3z-1?

OpenStudy (anonymous):

*simplest ?

OpenStudy (anonymous):

x^4y^6z^4

OpenStudy (anonymous):

yes, that is as simplest it could get

OpenStudy (anonymous):

He's wrong

OpenStudy (anonymous):

Wait he is right

OpenStudy (anonymous):

My mistake

OpenStudy (anonymous):

x5y3z3/xy-3z-1 x^5 y^3 z^3/ x y^-3 z^-1 = x^5 y^3 z^3 * y^3 z / x = x^4 y^6 z^4